At midnight a train left for Denver bound for Chicago, and another train left Chicago bound for Denver, both traveling at consatent speeds on adjacent tracks. The first train took 12 hours to complete the trip, and the second train took 16 hours to complete the trip. At what time did they passeach other?
Answer by checkley77(12569)

AVERAGE TIME=(12+16)/2=14 HOURS.
THUS THE HALF WAY POINT WOULD BE 14/2=7 HOURS (7:00 AM) INTO THE TRIP THEY WOULD PASS EACH OTHER.